# Varco Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models

### v1

@@ 7 Conclusion
We proposed Varco Arena, a reference-free LLM benchmarking approach for measuring generative capabilities of LLMs. Varco Arena deploys a tournament-style approach with direct pairwise comparisons, eliminating the need for reference outputs. Our simulation experiments demonstrate that this tournament approach achieves aligned LLM rankings better than current practice of reference use, under unbiased judging conditions. We also report the empirical effectiveness of Varco Arena using gpt-4o[-mini] as a judge, which reaffirms the efficacy of our approach in practice by showing improved correlation with Chatbot Arena leaderboard rankings. Through extensive experimentation and rigorous statistical analysis, we strongly suggest that our reference-free tournament approach, Varco Arena, provides a superior solution for LLM benchmarking, requiring fewer computational resources.
@@ A.1 Machine Requirements for Experiments
All the experiments we performed are conducted by querying API and post-processing those, which does not require GPU machine locally. Experiments could be run on personal desktops. The lowest specification of the machine we deployed had i5-8400 CPU, 16 GiB RAM.
Figure 4: Elo ratings of the model with gpt-4o judge on the full set of Arena-Hard-Auto (Li et al. 2024) prompts. Varco Arena result (bootstrapped median over 1000 samples of 500 trials) is in blue, plotted alongside the ratings from the ground truth leaderboard in red (Chatbot Arena, Hard prompts category). Error bars are 95% confidence intervals.
A.2 Elo ratings from Varco Arena compared to Human Annotations
@@
A.3 Binary search vs. Win rate over reference
@@
Converting Elo Table back to Win Rate
@@
Table 3: Binary search vs. Anchored comparison: Mean rank deviation ($|\Delta_{\text{rank}}|$) from ground-truth leaderboard. Result of binary search placement and anchored comparison insert by gpt-4o[-mini] judge are provided with bootstrapped 95% confidence interval (500 trials, 1000 samples, $|X|$=500, Arena-Hard-Auto (Li et al. 2024)).
A.4 Separability In terms of Confidence Interval
@@
A.5 Judge configuration
Evaluation Prompt
We use the prompt from LLMBar. The prompt depicted in Figure 1. We added 4 questions for criteria of our own to Metrics.txt prompt of (Zeng et al. 2024). You can refer to the original prompt in LLMBar github.
Decoding Parameters
@@ A.5 Judge configuration
Listing 1: LLMBar prompt (Metrics.txt)
1PROMPTS = [
2 {
3 role”: system”,
@@
4 content”: ”””You are a helpful assistant in evaluating the quality of the outputs for a given instruction. Your goal is to select the best output for the given instruction.”””,
5 },
6 {
7 role”: user”,
@@
8 content”: ”””Select the Output (a) or Output (b) that is better for the given instruction. The two outputs are generated by two different AI chatbots respectively.
9
10Here are some rules of the evaluation:
@@
11(1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc.
@@
12(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
@@
13(3) You should avoid any potential bias and your judgment should be as objective as possible. For example, the order in which the outputs were presented should NOT affect your judgment, as Output (a) and Output (b) are **equally likely** to be the better.
14
15Do NOT provide any explanation for your choice.
16Do NOT say both / neither are good.
17You should answer using ONLY Output (a)” or Output (b)”. Do NOT output any other words.
18
19# Instruction:
20{instruction}
21
22# Output (a):
23{response_a}
24
25# Output (b):
26{response_b}
27
28# Questions about Outputs:
@@
29Here are at most three questions about the outputs, which are presented from most important to least important. You can do the evaluation based on thinking about all the questions.
30* Does the output well satisfy the intent of the user request?
31* If applicable, is the output well-grounded in the given context information?
@@
32* Does the output itself satisfy the requirements of good writing in terms of:
33 1) Coherence
34 2) Logicality
35 3) Plausibility
36 4) Interestingness
37
38
39# Which is better, Output (a) or Output (b)? Your response should be either Output (a)” or Output (b)”:”””,
40 },
41]
Appendix B Reproducibility Checklist
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes)
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes)
  • Provides well marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper (yes)
  • Does this paper make theoretical contributions? (yes)
  • All assumptions and restrictions are stated clearly and formally. (yes)
  • All novel claims are stated formally (e.g., in theorem statements). (partial)
  • Proofs of all novel claims are included. (partial)
  • Proof sketches or intuitions are given for complex and/or novel results. (yes)
  • Appropriate citations to theoretical tools used are given. (yes)
  • All theoretical claims are demonstrated empirically to hold. (yes)
  • All experimental code used to eliminate or disprove claims is included. (no)
  • Does this paper rely on one or more datasets? (yes)
  • A motivation is given for why the experiments are conducted on the selected datasets (yes)
  • All novel datasets introduced in this paper are included in a data appendix. (NA)
  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes)
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes)
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes)
  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. (NA)
  • Does this paper include computational experiments? (yes)
  • Any code required for pre-processing data is included in the appendix. (no)
  • All source code required for conducting and analyzing the experiments is included in a code appendix. (no)
  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes)
  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (partial)
  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (yes)
  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. (partial)
  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes)
  • This paper states the number of algorithm runs used to compute each reported result. (yes)
  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (yes)
  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes)
  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes)
  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes)

### v2

@@ 7 Conclusion
We introduced VARCO Arena, a reference-free LLM benchmarking approach that employs a tournament-style framework with direct pairwise comparisons to evaluate the generative capabilities of LLMs. VARCO Arena offers a cost-effective, scalable, and adaptable solution for benchmarking response quality. Our simulation study and empirical experiments demonstrate that VARCO Arena consistently achieves higher rank reliability compared to current reference-based methods, as evidenced by stronger correlations with human-established Elo ratings. Given its robust performance and flexibility, we believe VARCO Arena can serve as a reliable automated tool for model selection and ranking across diverse and evolving use cases. Future work will explore broader applications, such as benchmarking LLMs in multi-modal settings (e.g., those incorporating visual or audio inputs and outputs).
Limitations
Although we tested robustness of the tournaments performed by VARCO Arena, it still has added factor for randomness such as match bracket initialization which does not applies to reference-based method. Rooms for improvement exist for more informative match-making algorithm that would achieve better ranking than single-elimination tournaments within same or less number of matches.
@@ A.1 Machine Requirements for Experiments
Except the part we inferenced open-weight models such as Llama, Qwen and Gemma, our experiments are mostly do not require GPU usage. Inference are done on one A100 GPU, but T4 would be enough for reproducing our experiments. Otherwise, our experiments require querying API and post-processing those with CPU. Experiments could be run on personal desktops. The lowest specification of the machine we deployed had i5-8400 CPU, 16 GiB RAM.
A.2 Assuring Statistical Significance of the Results within Budget for proprietary models
Figure 4: Elo ratings of the model with gpt-4o judge on the full set of Arena-Hard-Auto Li et al. (2024) prompts. VARCO Arena result (bootstrapped median over 1000 samples of 500 trials) is in blue, plotted alongside the ratings from the ground truth leaderboard in red (Chatbot Arena, Hard prompts category). Error bars are 95% confidence intervals.
A.3 Elo ratings from VARCO Arena compared to Human Annotations
@@
A.4 Binary search vs. Win rate over reference
@@ A.4.1 Binary Search
A.4.2 Comparing to the most performant Model so far: Converting Elo Table back to Win Rate
@@
Table 5: Binary search vs. Anchored comparison: Mean rank deviation ($|\Delta_{\text{rank}}|$) from ground-truth leaderboard. Result of binary search placement and anchored comparison insert by gpt-4o[-mini] judge are provided with bootstrapped 95% confidence interval (500 trials, 1000 samples, $|X|$=500, Arena-Hard-Auto Li et al. (2024)).
A.5 Separability In terms of Confidence Interval
@@
A.6 Judge configuration
A.6.1 Evaluation Prompt
We use the prompt from LLMBar. The prompt depicted in Figure A.6.2. We added 4 questions for criteria of our own to Metrics.txt prompt of Zeng et al. (2024). You can refer to the original prompt in LLMBar github.
A.6.2 Decoding Parameters
@@
PROMPTS = [ # metrics.txt from LLMBar
@@
"role": "system", "content": "You are a helpful assistant in evaluating the quality of the outputs for a given instruction. Your goal is to select the best output for the given instruction.", ,
@@
"role": "user", "content": """Select the Output (a) or Output (b) that is better for the given instruction. The two outputs are generated by two different AI chatbots respectively.
Here are some rules of the evaluation:
@@
(1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc.
@@
(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
@@
(3) You should avoid any potential bias and your judgment should be as objective as possible. For example, the order in which the outputs were presented should NOT affect your judgment, as Output (a) and Output (b) are **equally likely** to be the better.
Do NOT provide any explanation for your choice.
Do NOT say both / neither are good.
You should answer using ONLY "Output (a)" or "Output (b)". Do NOT output any other words.
# Instruction:
instruction
# Output (a):
response_a
# Output (b):
response_b
# Questions about Outputs:
@@
Here are at most three questions about the outputs, which are presented from most important to least important. You can do the evaluation based on thinking about all the questions.
- Does the output well satisfy the intent of the user request?
- If applicable, is the output well-grounded in the given context information?
@@
- Does the output itself satisfy the requirements of good writing in terms of:
1) Coherence
2) Logicality
3) Plausibility
4) Interestingness
# Which is better, Output (a) or Output (b)? Your response should be either "Output (a)" or "Output (b)":""",